OpenStudy (anonymous):

translate each problem into an equation. the legnth of a rectangle is one unit more than its width. if the area is 30 square units, find the dimensions of the rectangle.

OpenStudy (btaylor):

L = 1+W L x W = 30 W(1+W)=30 W^2 + W-30 = 0 (W-5)(W+6) = 0 W = 5 (width cannot be negative) L = 1+W = 1+5 = 6

OpenStudy (anonymous):

thanks..could you help me with this one.. a triangle has two equal sides and a third side that is 15cm long. if the perimeter is 50 cm, how long is each of the two equal sides

OpenStudy (btaylor):

a + b + c = P a=b --> a + a + 15 = 50 2a= 35 a = 17.5 --> b = 17.5

OpenStudy (anonymous):

thanks!